# Restricted-Geometry Quantum Models Beyond Atoms: Application to NSDI in Diatomic Systems

### v1

@@ Abstract
We present a (1+1)-dimensional quantum model designed to describe nonsequential double ionization (NSDI) in homonuclear diatomic molecules exposed to strong linearly polarized laser fields. Extending the restricted-geometry framework previously developed for atomic systems, our approach captures key features of NSDI, including the characteristic knee structure in double ionization yields. Despite its simplifying assumptions, the model shows good agreement with experimental data and proves particularly suitable for systems with -type orbital symmetry. It offers a computationally efficient tool for exploring multi-electron dynamics in molecular systems.
@@ I Introduction
Building on this foundational model, Eckhardt and Sacha proposed a new perspective on non-sequential double ionization (NSDI) of atoms [5, 6, 7, 8, 9, 10], which led to the introduction of a novel restricted-geometry quantum model [11, 12]. This model successfully reproduced and explained key experimental observations, such as the knee structure in the ionization yield curve [13, 14, 15, 16], the double-hump structure in the ion momentum distribution [17, 18], and characteristic structures in two-electron momentum distributions [19, 20, 21], while also enabling the analysis of NSDI across a broad range of laser field parameters [22, 23, 24, 25, 26, 27] .
@@
It is evident that models developed for atomic systems fall short in capturing the molecular-specific effects. To address this, theoretical approaches must be adapted or extended to incorporate parameters such as internuclear separation and orientation-dependent ionization pathways. In this work, we aim to extend the restricted-geometry quantum framework to diatomic molecules and assess its ability to capture both universal and molecule-specific features of strong-field ionization. Our findings aim to contribute to a more unified understanding of multi-electron dynamics in atoms and molecules under intense laser fields.
@@ II.2 Restricted-Geometry Hamiltonian
Sacha and Eckhardt suggested in [7, 6] that correlated double ionization requires the electrons to escape simultaneously, crossing a pair of saddles formed by the interplay between nuclear attraction, interaction with the electric field, and electron–electron repulsion. Naturally, in the molecular case, the three two-dimensional potentials (8), (9), and (10) exhibit saddles located at different positions and with different properties, which may affect the ionization signals.
@@
For the configurations represented by (Eq. (8), Fig. 1(a)) and (Eq. (9), Fig. 1(b)), we apply the coordinate transformation , , and for [10], in Eqs. (5) and (7). In the third configuration, represented by (Eq. (10), Fig. 1(c)), the roles of and are interchanged. For notational compactness, we define the set , and in the expressions below, indicates that the potential is computed as a sum over all values in this set.
The electron–core interaction terms for the three configurations are then given by
@@ II.3 Model Assumptions and Limitations
Since our aim in this work is to explore the predictive power and limitations of the restricted-geometry quantum framework as applied to NSDI in diatomic molecules, a few comments are in order before we proceed to the numerical implementation and discussion of the results.
First, we note that the starting point of our approach is a hydrogen-like Hamiltonian. At this level, the internal structure of the target (atom or molecule) is not explicitly included, similar to the SAE approximation commonly used in the description of HHG or ATI. Regardless of the atomic species, SAE assumes that only one electron interacts with the field, while the influence of the remaining electrons is incorporated into an effective potential. In our case, we include two active electrons, and the only parameter in the Hamiltonian that reflects the molecular species is the internuclear distance . In this sense, the model describes hydrogen-molecule-like systems. However, in the numerical implementation (see next section), we also introduce a parameter that allows us to set the ionization potential, thereby tailoring the model to a specific molecular species.
An open question remains: to what extent can molecular symmetries, such as the distinction between and orbitals, be mapped onto the initial wave function within the restricted-geometry quantum framework? Similar considerations have been made in the context of differences between the ground and excited states of helium [23], as well as in studies of triple ionization [29, 30, 31, 32, 33, 34]. This issue may serve as a fruitful direction for future research, particularly in the context of extending the model to more complex molecular systems.
@@ IV Results
In the following subsections, we compare ionization yield curves for different molecular alignments and species, and present results for momentum distributions. We emphasize once again that the analysis of molecular alignment is limited to two extreme orientations, parallel and perpendicular, while the molecular species differ only in their internuclear distance and ionization energy. In the diagrams, ionization yields are color-coded for clarity and consistency across comparisons: red for N2, blue for O2, and yellow for S2.
@@ IV.1 Ionization Yields
To further compare the studied molecular species, we apply the classical scaling described in the previous section. As an example, we first show a comparison of ionization yields for the perpendicular orientation () between N2 and O2 in Fig. 6, and for the parallel orientation () between N2 and S2 in Fig. 7. In both cases, a remarkably good agreement in the global trend is observed, highlighting once again the model’s limited ability to distinguish between molecular species.
@@
To test this behavior, we repeated our simulations using a trapezoidal pulse shape and verified that the extrema become more pronounced with increasing pulse duration. This is illustrated in Fig. 9 for the case of S2 and potential . For longer pulses, the peaks can be associated with resonances among eigenstates of the periodically driven system, i.e., its Floquet states. Similar phenomena have been observed in microwave ionization of atomic systems [89, 90], where they were directly linked to Floquet resonances. Although we do not explicitly construct Floquet states in the present model, the stabilization of the ionization yield with increasing pulse duration supports the interpretation in terms of resonance effects.

### v2

# Restricted-Geometry Quantum Models Beyond Atoms: Application of the Eckhardt-Sacha approach to NSDI in Diatomic Systems

@@ Abstract
We present a (1+1)-dimensional quantum model designed to describe nonsequential double ionization (NSDI) in homonuclear diatomic molecules exposed to strong linearly polarized laser fields. Extending the restricted-geometry framework previously developed for atomic systems by Eckhardt and Sacha, our approach captures key features of NSDI, including the characteristic knee structure in double ionization yields. Despite its simplifying assumptions, the model shows good agreement with experimental data and proves particularly suitable for systems with -type orbital symmetry. It offers a computationally efficient tool for exploring multi-electron dynamics in molecular systems.
@@ I Introduction
These developments were accompanied by the extensive use of another approach that greatly simplifies calculations and provides intuitive explanations of the studied phenomena [5, 6, 7, 8, 9, 10, 11, 12, 13]. This approach is based on the observation that, in a linearly polarized field, the dominant influence on electron dynamics occurs along the polarization axis. Consequently, the generally three-dimensional problem for a single electron can be reduced to a one-dimensional model that captures the essential physics of the analyzed effect. For non-sequential double ionization (NSDI), this simplification results in a -dimensional model, with one coordinate along the laser polarization axis for each electron [10, 11, 12, 14, 13].
Building on these simplified and restricted-geometry approaches, Eckhardt and Sacha introduced a new perspective on NSDI of atoms [15, 16, 17, 18, 19, 20], which culminated in the development of a novel restricted-geometry quantum model [21, 22]. The key idea of this approach is to constrain the escaping electrons to evolve within a symmetric subspace that naturally reflects the correlated two-electron motion characteristic of NSDI, thereby isolating the physically relevant degrees of freedom while retaining the essential recollision-driven dynamics. This model successfully reproduced and explained key experimental observations, such as the knee structure in the ionization yield curve [23, 24, 25, 26], the double-hump structure in the ion momentum distribution [27, 28], and characteristic structures in two-electron momentum distributions [29, 30, 31], while enabling systematic analysis of NSDI across a broad range of laser-field parameters [32, 33, 34, 35, 36, 37, 38].
@@
It is evident that models developed for atomic systems fall short in capturing molecular-specific effects. To address this, theoretical approaches must be adapted or extended to incorporate parameters such as internuclear separation and orientation-dependent ionization pathways, just to mention the simplest ones. Yet, moving from atomic to molecular models inevitably increases computational complexity and the numerical cost of full-dimensional simulations. This raises a fundamental and still open question: to what extent can simplified restricted-geometry models be reliably applied to molecular systems? Before introducing our model, we note that simplified approaches to molecular NSDI also necessarily rely on assumptions such as fixed nuclei and the use of smoothed Coulomb interactions; these aspects, together with their implications, are discussed in Sec. II.
As long as one considers models in which linear polarization of the laser field defines a privileged direction of electron motion, the geometric reduction is straightforward and follows the same logic as in the atomic case. However, when the restricted-geometry model originally derived by Eckhardt and Sacha for atoms [15, 16, 17, 18, 19, 20] is to be applied to molecular systems, its validity and implementation require careful justification. A classical extension of this approach to molecules was proposed in [81, 82], but a quantum counterpart has not been systematically developed or evaluated.
In this work, we extend the restricted-geometry quantum framework within the Eckhardt–Sacha approach to diatomic molecules and assess its ability to capture both universal and molecule-specific features of strong-field ionization. Our findings aim to clarify the applicability limits of this method and to identify which molecular NSDI mechanisms can be reproduced within such reduced-dimensionality models. Additionally, the approach offers a computationally efficient tool for studying multi-electron dynamics in systems where full-dimensional quantum calculations remain prohibitively expensive.
@@ II.2 Restricted-Geometry Hamiltonian
[…]
@@
[…]
@@ II.3 Model Assumptions and Limitations
Since our more specific goal in this work is to explore the predictive power and limitations of the restricted-geometry quantum framework within Eckhard-Sacha approach as applied to NSDI in diatomic molecules, a few comments are in order before we proceed to the numerical implementation and discussion of the results.
First, we note that the starting point of our approach is a hydrogen-like Hamiltonian. At this level, the internal structure of the target (atom or molecule) is not explicitly included, similar to the SAE approximation commonly used in the description of HHG or ATI. Regardless of the atomic species, SAE assumes that only one electron interacts with the field, while the influence of the remaining electrons is incorporated into an effective potential. In our case, we include two active electrons, and the only parameter in the Hamiltonian that reflects the molecular species is the internuclear distance . In this sense, the model describes hydrogen-molecule-like systems. However, in the numerical implementation (see sec. III), we also introduce a parameter that allows us to set the ionization potential, thereby tailoring the model to a specific molecular species.
In the following, we start with the two-electron ground state function that has spacial part symmetric under particle exchange, i.e. spin part corresponds to a singlet - orbital. An open question remains: to what extent can molecular symmetries, such as the distinction between and orbitals, be mapped onto the initial wave function within the any restricted-geometry quantum framework? Similar considerations have been made in the context of differences between the ground and excited states of helium [33], as well as in studies of triple ionization [40, 41, 42, 43, 44, 45]. This issue may serve as a fruitful direction for future research, particularly in the context of extending the model to more complex molecular systems.
@@ III Numerical approach
The usefulness of this classical scaling lies in its ability to compensate for the differences in ionization energies between the model potentials. Since the ground-state energy sets the characteristic scale of electron motion, molecules with different respond most similarly when the driving field parameters are rescaled according to Eqs. (21)–(22). In practice, this scaling aligns the effective barrier heights and excursion amplitudes experienced by the electrons, allowing ionization yields of systems with different binding energies to be compared on an approximately equal footing. Although the procedure is not exact, since it combines quantum-mechanical energies with a classical scaling, it captures the dominant trends and substantially improves the interpretability of cross-molecular ionization-yield comparisons.
@@ IV Results
In the following subsections, we compare ionization yield curves for different molecular alignments and species, and present results for momentum distributions. We emphasize once again that the analysis of molecular alignment is limited to two extreme orientations, parallel and perpendicular, while the molecular species differ only in their internuclear distance and ionization energy.
@@ IV.1 Ionization Yields
To further compare the studied molecular species, we apply the classical scaling described in the previous section. As an example, we first show a comparison of ionization yields for the perpendicular orientation () between N2 and O2 in Fig. 6, and for the parallel orientation () between N2 and S2 in Fig. 7. In both cases, a remarkably good agreement in the global trend is observed, highlighting once again the model’s limited ability to distinguish between molecular species (i.e., the molecular species are differentiated only by a change in the internuclear distance , not by their ground‑state symmetry).
@@
To test this behavior, we repeated our simulations using a trapezoidal pulse shape and verified that the extrema become more pronounced with increasing pulse duration. This is illustrated in Fig. 9 for the case of S2 and potential . For longer pulses, the peaks can be associated with resonances among eigenstates of the periodically driven system, i.e., its Floquet states. Following the approach of Panfiliet al. [95], we estimated the positions of multiphoton resonances in our model by including both the dynamical Stark shift and the ponderomotive shift of the bound-state energies. These estimates, as applied to results for S2 shown in Fig. 9, indicate that 3-, 4-, and 5-photon resonances between the ground state and the first or second excited state should occur within the same range of field strengths where the local extrema in the double-ionization yield are observed. In this picture, the resonances enhance population transfer into Stark- and field-dressed excited states that act as intermediate channels for double ionization, thereby producing the characteristic peaks. Similar phenomena have been observed in microwave ionization of atomic systems [96, 97], where they were directly linked to Floquet resonances. Although we do not explicitly construct Floquet states in the present model, the stabilization of the ionization yield with increasing pulse duration supports the interpretation in terms of resonance effects.
@@ V Conclusions
To summarize the physical content of the model, we emphasize which experimental observations it can and cannot reproduce. On the positive side, the model correctly captures (i) the existence and approximate position of the knee structure in the DI yield, (ii) the global dependence of SI and DI yields on the field strength, (iii) resonance-induced non‑monotonic features such as the local extrema traced to multiphoton and Floquet-type resonances, and (iv) characteristic signatures in two-electron and ion momentum distributions, including asymmetric energy sharing and flat-top ion momentum profiles, consistent with experimental observations. However, the model cannot reproduce (i) the experimentally observed dependence of SI and DI yields on molecular orientation, (ii) the suppression of DI in molecules with -type valence orbitals, such as O2, (iii) effects associated with multielectron orbital symmetry and configuration mixing, and (iv) orientation averaging or nuclear-motion–induced modifications of ionization channels. These limitations stem directly from the restricted geometry, the absence of explicit orbital symmetry, and the hydrogenic character of the underlying effective potential. A clear identification of which observables fall within the model’s predictive scope helps guide the interpretation of agreements and deviations with experimental data.
@@ Acknowledgements.
- Javanainen et al. [1988] J. Javanainen, J. H. Eberly, and Q. Su, Numerical simulations of multiphoton ionization and above-threshold electron spectra, Phys. Rev. A 38, 3430 (1988).
- Eberly et al. [1989] J. H. Eberly, Q. Su, and J. Javanainen, Nonlinear light scattering accompanying multiphoton ionization, Phys. Rev. Lett. 62, 881 (1989).
- Eberly [1990] J. H. Eberly, Scale variation in a one-dimensional model of an atom interacting with a strong laser field, Phys. Rev. A 42, 5750 (1990).
- Su et al. [1990] Q. Su, J. H. Eberly, and J. Javanainen, Dynamics of atomic ionization suppression and electron localization in an intense high-frequency radiation field, Phys. Rev. Lett. 64, 862 (1990).
- Pindzola et al. [1991] M. S. Pindzola, D. C. Griffin, and C. Bottcher, Validity of time-dependent hartree-fock theory for the multiphoton ionization of atoms, Phys. Rev. Lett. 66, 2305 (1991).
- Becker et al. [2012] W. Becker, X. Liu, P. J. Ho, and J. H. Eberly, Theories of photoelectron correlation in laser-driven multiple atomic ionization, Rev. Mod. Phys. 84, 1011 (2012).
@@
- Liu et al. [2025] H. Liu, H. Zhang, X. Wang, and J. Yuan, Atomic double ionization with quantum light, Phys. Rev. Lett. 134, 123202 (2025).